The variance of 20 observations is 5. If each observation is multiplied by 2, then find the new
variance of the resulting observations.

Grade:upto college level

2 Answers

Latika Leekha
9 years ago
Hello student,
Let the given observations be x1, x2, …. x20. and xM be their mean.
Given that n = 20 and variance is 5.
So, \frac{1}{20}\sum_{i =1}^{20}(x-x_{M})^{2} = 5
\sum_{i=1}^{20}(x-x_{M})^{2}=100 ..... (1)
If each observation is multiplied by 2, then the new observations are yi = 2xi, i = 1, 2, … 20.
So, xi = yi/2.
Hence, y_{M}= \frac{1}{20}\sum_{i=1}^{20}y_{i}
= \frac{1}{20}\sum_{i=1}^{20}2x_{i}
= 2. 1/20 . xM
This gives us xM = ½ yM
Substituting the values of xi and xM in (1) we have
\sum_{i=1}^{20}(\frac{1}{2}y_{i}-\frac{1}{2}y_{M})^{2}=100
\sum_{i=1}^{20}\frac{1}{4}(y_{i}-y_{M})^{2}=100
\sum_{i=1}^{20}\(y_{i}-y_{M})^{2}=400
Hence, the variance of the resulting observations is = 1/20 . 400
= 20.
= 22.5
ankit singh
3 years ago

Answer

If each observation is multiplied by 2, then mean(μ) will get doubled.

Now, variance =σ2=n(Xμ)2
Now here both X and μ will be doubled, so the variance will become four times.
Thus, the new variance will be 4 times the older variance which is 4×5=20.